# Universal $L_{2}$-approximation using median lattice algorithms

### v1

@@ Abstract
We study the problem of multivariate $L_{2}$-approximation of functions in a weighted Korobov space using a median lattice-based algorithm recently proposed by the authors. In the original work, the algorithm requires knowledge of the smoothness and weights of the Korobov space to construct the hyperbolic cross index set, where each coefficient is estimated via the median of approximations obtained from randomly shifted, randomly chosen rank-1 lattice rules. In this paper, we introduce a universal median lattice-based algorithm, which eliminates the need for any prior information on smoothness and weights. Although the tractability property of the algorithm slightly deteriorates, we prove that, for individual functions in the Korobov space with arbitrary smoothness and (downward-closed) weights, it achieves an $L_{2}$-approximation error arbitrarily close to the optimal rate with respect to the number of function evaluations.
@@ 1 Introduction
The rest of the paper is structured as follows. In Section 2, we introduce the necessary preliminaries and notation, including the weighted Korobov space, the $L_{2}$-approximation problem, and rank-1 lattice rules. Section 3 presents the universal median lattice-based algorithm and states the main probabilistic $L_{2}$-error bound (Theorem 3.8), along with a discussion of its tractability properties. In Section 4, we provide a detailed analysis of the three terms in the error decomposition ($\clubsuit$3.2), establishing probabilistic bounds that, when combined, yield the main result. We conclude this paper with a discussion in Section 5.
@@ 3.3 Tractability analysis
First, let us consider the quantity $\bar{r}_{\mathrm{sum}}$ defined in (17). By Lemma 3.4, for any $\lambda\in(1/(2\alpha),1)$ we obtain
@@
for any $\lambda\in(1/(2\alpha),1)$, where $C_{1},C_{2},C_{1,\alpha,\lambda},C_{2,\alpha,\lambda}$ are constants independent of $N$, $R$, and $d$, but possibly depending on $\alpha$ and $\lambda$, as indicated by their subscripts.
@@
If, for some $\lambda\in(1/(2\alpha),1)$ and some $q>0$, we have $B_{\lambda,q}<\infty$ and $\widetilde{B}_{q}<\infty$, then the error bound in (6) depends at most polynomially on $d$, and an error of at most $\varepsilon>0$ can be achieved with $N$ depending at most polynomially on $d$ and $\varepsilon$, with high probability. Moreover, the number $R$ of independent rank-1 lattices required for this result depends only logarithmically on $N$ and linearly on $d$ by (5). This situation is referred to as polynomial tractability in the context of Information-Based Complexity (see, e.g., [21]).
@@
for some $\lambda\in(1/(2\alpha),1)$. This assumption implies that, for any fixed $\delta>0$, there exists $d_{\delta}$ such that
@@
$\leq\frac{1}{d^{q}}\exp\left(2(A_{\lambda}+\delta)\zeta(2\alpha\lambda)\log d)\right)=d^{2(A_{\lambda}+\delta)\zeta(2\alpha\lambda)-q}.$
@@ 4 Error analysis: bounds on the error decomposition
The next lemma illustrates how the median trick reduces the failure probability. It comes from [16, Proposition 2.2], which is a minor modification of [20, Proposition 2.1].
@@
$\bar{S}(\boldsymbol{h}):=4\left(\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N-1}+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\right)^{1/2},$
and
@@
where $N_{*}$ and $r_{\mathrm{sum}}(\boldsymbol{h})$ are as defined in (2) and Lemma 4.8, respectively.
@@
$\delta^{\prime\prime}=4\left(\frac{31}{15}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}\right)^{1/2}.$
@@
$\Pr\left(\operatorname*{median}_{r\in\{1{:}R\}}\left|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})\right|^{2}>\delta^{\prime\prime 2}\right)\leq\frac{1}{2}\,2^{-R/2}.$
@@
If $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)\setminus K^{\prime}_{N}$, then by construction
$F(\boldsymbol{h})\leq F(\boldsymbol{h}^{\prime}_{N})\leq\frac{16}{15}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)},$
where the second inequality follows from Lemma 4.11. Substituting this bound into the definition of $\delta^{\prime}(\boldsymbol{h})$ in Lemma 4.2 gives $\delta^{\prime}(\boldsymbol{h})^{2}\leq\delta^{\prime\prime 2}$. The claimed probability estimate then follows immediately from Lemmas 4.2 and 4.3. ∎
@@
where both $\bar{S}(\boldsymbol{h})$ and $\bar{r}_{\mathrm{sum}}$ are as defined in Lemma 4.9, then
@@
    $\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\delta^{\prime\prime 2}.$
@@
$\frac{F(\boldsymbol{h}^{\prime})}{2}<\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}^{\prime})|^{2}\leq\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\delta^{\prime\prime 2},$
@@
$|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}\leq F(\boldsymbol{h}^{\prime})\leq 2\delta^{\prime\prime 2}.$
@@
$\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}\leq|K^{\prime}_{N}\setminus K_{N}|\cdot 2\delta^{\prime\prime 2}+\max\left(960N^{-2\alpha}_{*},\sqrt{240\bar{r}_{\mathrm{sum}}}\right)\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}\bar{S}^{2}(\boldsymbol{h}^{\prime}).$
@@ 5 Discussion
In this work, we have extended the median lattice-based algorithm from [22] to the setting where no prior knowledge of the smoothness $\alpha$ and the weight parameters $\boldsymbol{\gamma}$ of the Korobov space is available. A major limitation of our method, as pointed out in Remark 3.3, lies in Step 2 of Algorithm 3.1, where we need to order $\mathrm{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}$ for $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$ and select the $N$ largest frequencies $K_{N}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\}$. A naive implementation requires computing $\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})$ for all $r\in\{1{:}R\}$ and for every $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$, whose cardinality grows exponentially with $d$. It is therefore natural to ask whether $K_{N}$, or an adequate substitute, can be found in computation time polynomial in $d$. For example, a mild assumption on $\alpha$ and $\boldsymbol{\gamma}$ may suffice to identify a polynomial-size subset $\mathcal{A}^{*}\subseteq\mathcal{A}_{d}(N/2)$ that contains $K_{N}$. In the extreme case where $\alpha$ and $\boldsymbol{\gamma}$ are fully known, $K_{N}$ can be replaced by the set $\mathcal{A}_{d}(N_{*})$ defined in [22] while still achieving a near $O(M^{-\alpha})$ convergence rate. Another possible approach is to estimate $\alpha$ and $\boldsymbol{\gamma}$ empirically by computing $\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})$ over a small test set for $r\in\{1{:}R\}$. We leave this question for future research.

### v2

@@ Abstract
We study the problem of multivariate $L_{2}$-approximation of functions in a weighted Korobov space using a median lattice-based algorithm recently proposed by the authors. In the original work, the algorithm requires knowledge of the smoothness and weights of the Korobov space to construct the hyperbolic cross index set, where each coefficient is estimated via the median of approximations obtained from randomly shifted, randomly chosen rank-1 lattice rules. In this paper, we introduce a universal median lattice-based algorithm, which eliminates the need for any prior information on smoothness and weights. Although the tractability property of the algorithm slightly deteriorates, we prove that, for individual functions in the Korobov space with arbitrary smoothness and (downward-closed) weights, it achieves an $L_{2}$-approximation error arbitrarily close to the optimal rate with respect to the number of function evaluations. Numerical experiments are conducted to support our theoretical claim.
00footnotetext: Email addresses: zep002@zju.edu.cn (Zexin Pan) ; goda@frcer.t.u-tokyo.ac.jp (Takashi Goda); peter.kritzer@oeaw.ac.at (Peter Kritzer)
@@ 1 Introduction
The rest of the paper is structured as follows. In Section 2, we introduce the necessary preliminaries and notation, including the weighted Korobov space, the $L_{2}$-approximation problem, and rank-1 lattice rules. Section 3 presents the universal median lattice-based algorithm and states the main probabilistic $L_{2}$-error bound (Theorem 3.8), along with a discussion of its tractability properties. In Section 4, we provide a detailed analysis of the three terms in the error decomposition ($\clubsuit$3.2), establishing probabilistic bounds that, when combined, yield the main result. In Section 5, we conduct numerical experimetns to support our theoretical claim. We conclude this paper with a discussion in Section 6.
@@ 3.3 Tractability analysis
First, let us consider the quantity $\bar{r}_{\mathrm{sum}}$ defined in (5). By Lemma 3.4, for any $\lambda\in(1/(2\alpha),2)$ we obtain
@@
for any $\lambda\in(1/(2\alpha),2)$, where $C_{1},C_{2},C_{1,\alpha,\lambda},C_{2,\alpha,\lambda}$ are constants independent of $N$, $R$, and $d$, but possibly depending on $\alpha$ and $\lambda$, as indicated by their subscripts.
@@
If, for some $\lambda\in(1/(2\alpha),2)$ and some $q>0$, we have $B_{\lambda,q}<\infty$ and $\widetilde{B}_{q}<\infty$, then the error bound in (7) depends at most polynomially on $d$, and an error of at most $\varepsilon>0$ can be achieved with $N$ depending at most polynomially on $d$ and $\varepsilon$, with high probability. Moreover, the number $R$ of independent rank-1 lattices required for this result depends only logarithmically on $N$ and linearly on $d$ by (6). This situation is referred to as polynomial tractability in the context of Information-Based Complexity (see, e.g., [22]).
@@
for some $\lambda\in(1/(2\alpha),2)$. This assumption implies that, for any fixed $\delta>0$, there exists $d_{\delta}$ such that
@@
$\leq\frac{1}{d^{q}}\exp\left(2(A_{\lambda}+\delta)\zeta(2\alpha\lambda)\log d\right)=d^{2(A_{\lambda}+\delta)\zeta(2\alpha\lambda)-q}.$
@@ 4 Error analysis: bounds on the error decomposition
The next lemma illustrates how the median trick reduces the failure probability. It comes from [17, Proposition 2.2], which is a minor modification of [21, Proposition 2.1].
@@
$\bar{S}(\boldsymbol{h}):=4\left(\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N-1}+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\right)^{1/2}.$ (18)
@@
$\bar{\delta}=4\left(\frac{31}{15}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}\right)^{1/2}.$
@@
$\Pr\left(\operatorname*{median}_{r\in\{1{:}R\}}\left|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})\right|^{2}>\bar{\delta}^{2}\right)\leq\frac{1}{2}\,2^{-R/2}.$
@@
If $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)\setminus K^{\prime}_{N}$, then by Lemma 4.11 and the definition of $K^{\prime}_{N}$
$F(\boldsymbol{h})\leq F(\boldsymbol{h}^{\prime}_{N})\leq\frac{16}{15}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}.$
The claimed probability estimate then follows immediately from Lemma 4.10. ∎
@@
where $\bar{S}(\boldsymbol{h})$ and $\bar{r}_{\mathrm{sum}}$ are as defined in (18) and (5), respectively, then
@@
    $\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\bar{\delta}^{2}.$
@@
$\frac{F(\boldsymbol{h}^{\prime})}{2}<\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}^{\prime})|^{2}\leq\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\bar{\delta}^{2},$
@@
$|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}\leq F(\boldsymbol{h}^{\prime})\leq 2\bar{\delta}^{2}.$
@@
$\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}\leq|K^{\prime}_{N}\setminus K_{N}|\cdot 2\bar{\delta}^{2}+\max\left(960N^{-2\alpha}_{*},\sqrt{240\bar{r}_{\mathrm{sum}}}\right)\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}\bar{S}^{2}(\boldsymbol{h}^{\prime}).$
@@
5 Numerical experiments
Here, we conduct numerical experiments to support our theoretical findings. The key feature of Algorithm 3.1 is its universality; it does not require any prior information on the smoothness $\alpha$ and the weights $\boldsymbol{\gamma}$ as an input, yet it automatically exploits the regularity of the target function. The focus of these experiments is to test this capability to exploit smoothness. To do this, we consider the simplest multivariate setting, i.e., the unweighted, bivariate case.
Following a previous paper by the same authors on median lattice-based algorithms [23], we test the following two periodic functions, which were also used in [3]:
$f_{1}(\boldsymbol{x})$ $=\prod_{j=1}^{d}\frac{121\sqrt{33}}{100}\max\left\{\frac{25}{121}-\left(x_{j}-\frac{1}{2}\right)^{2},0\right\},$
$f_{2}(\boldsymbol{x})$ $=\prod_{j=1}^{d}\left(x_{j}-\frac{1}{2}\right)^{2}\sin(2\pi x_{j}-\pi),$
with $d=2$. It is important to note that these functions belong to different smoothness classes: $f_{1}\in\mathcal{H}_{d,3/2-\epsilon,\boldsymbol{\gamma}}$ and $f_{2}\in\mathcal{H}_{d,5/2-\epsilon,\boldsymbol{\gamma}}$ for any arbitrarily small $\epsilon>0$. This means that the smoothness of $f_{1}$ is arbitrarily close to $3/2$ and that of $f_{2}$ is arbitrarily close to $5/2$. Consequently, based on Theorem 3.8 and Corollary 3.10, our universal median lattice-based algorithm is expected to achieve $L_{2}$ error rates arbitrarily close to $M^{-3/2}$ for $f_{1}$ and $M^{-5/2}$ for $f_{2}$, respectively, where $M=NR$ denotes the total number of function evaluations.
For our implementation of Algorithm 3.1, we set $R$ to the smallest odd integer that satisfies the condition in (6) with dimension $d=2$ and failure probability $\varepsilon=0.01$. This choice ensures that the error bound shown in Corollary 3.10 holds with a probability of at least $0.99$. For comparison, we also run the median lattice-based algorithm from [23]. Unlike our present method, this algorithm requires prior knowledge of the function’s smoothness and the weights to construct a proper index set. For a fair comparison, we use the same parameter settings and failure probability ($\varepsilon=0.01$) as reported in the original paper [23].
Figure 1 presents the experimental results, plotting the $L_{2}$ error of both algorithms against the total number of function evaluations, $M$, on a log-log scale. The left and right panels show the results for test functions $f_{1}$ and $f_{2}$, respectively. The $L_{2}$ error is calculated exactly by using the true Fourier coefficients and the $L_{2}$-norm of the functions, see [3, 23]. Each panel includes reference lines for the rates $M^{-\alpha/2}$, $M^{-3\alpha/4}$, and the optimal rate $M^{-\alpha}$, where the smoothness parameter is $\alpha=3/2$ for $f_{1}$ and $\alpha=5/2$ for $f_{2}$. For both functions, our universal algorithm significantly outperforms the non-universal baseline. This superior performance can be attributed to its ability to adaptively construct the index set $K_{N}$ based on the magnitude of the estimated Fourier coefficients. Nevertheless, the observed convergence rate for our algorithm is approximately $M^{-3\alpha/4}$ within the tested range of $M$, see below for an explanation for this phenomenon.
Figure 1: $L_{2}$-approximation error for the test functions $f_{1}$ (left) and $f_{2}$ (right) as a function of the total number of function evaluations $M$.
To evaluate how well our universal algorithm selects the index set $K_{N}$, we compute the normalized truncation error:
$\left(\frac{\sum_{\boldsymbol{k}\in\mathbb{Z}^{d}\setminus K_{N}}|\widehat{f}(\boldsymbol{k})|^{2}}{\|f\|^{2}_{L_{2}}}\right)^{1/2}=\left(1-\frac{\sum_{\boldsymbol{k}\in K_{N}}|\widehat{f}(\boldsymbol{k})|^{2}}{\|f\|^{2}_{L_{2}}}\right)^{1/2},$
where $K_{N}$ is the set of $N$ indices constructed in the second step of Algorithm 3.1, and $\widehat{f}(\boldsymbol{k})$ denotes the true $\boldsymbol{k}$-th Fourier coefficient. We benchmark this error against the theoretically optimal performance achievable with any set of $N$ indices. This benchmark is given by the oracle truncation error:
$\left(1-\frac{\sum_{\boldsymbol{k}\in K^{\mathrm{oracle}}_{N}}|\widehat{f}(\boldsymbol{k})|^{2}}{\|f\|^{2}_{L_{2}}}\right)^{1/2}.$
[…]
Figure 2 compares these two truncation errors, plotting them against the total number of function evaluations, $M$, on a log-log scale. Although the index set $K_{N}$ constructed by our algorithm is stochastic, the figure shows the result from a single realization. For both test functions, the curves from our algorithm and the oracle are nearly indistinguishable, which implies that the adaptive choice of $K_{N}$ is remarkably close to optimal. The convergence rate of this truncation error is again observed to be approximately $M^{-3\alpha/4}$ within the tested range. This finding explains the performance seen in Figure 1: the overall $L_{2}$ error is already limited by the convergence of the truncation error itself. We therefore conclude that the observed suboptimal convergence rate is a finite-sample effect, and achieving the theoretical rate of $M^{-\alpha}$ would likely require experiments in a larger-$M$ regime.
Figure 2: Normalized truncation error for the test functions $f_{1}$ (left) and $f_{2}$ (right). The error from the adaptive selection of the index set is compared to the optimal (oracle) error.
@@ 6 Discussion
This work extends the median lattice algorithm from [23] to the Korobov space setting with unknown smoothness $\alpha$ and weight parameters $\boldsymbol{\gamma}$. A primary computational bottleneck, as pointed out in Remark 3.3, lies in Step 2 of Algorithm 3.1, where we we must identify the $N$ largest values of $\mathrm{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}$ across the candidate set $\mathcal{A}_{d}(N/2)$, whose cardinality grows exponentially with the dimension $d$. A naive implementation of this step is therefore computationally prohibitive in high dimensions.
This motivates us to ask whether $K_{N}$, or a suitable substitute, can be found in computation time polynomial in $d$. A natural solution is to constrain the search to a manageable subset $\mathcal{A}^{*}\subseteq\mathcal{A}_{d}(N/2)$, designed within a given computational budget. The design of $\mathcal{A}^{*}$ can incorporate prior knowledge of $\alpha$ and $\boldsymbol{\gamma}$ if available. Indeed, with full knowledge, we recover the original algorithm from [23] by setting $\mathcal{A}^{*}$ to be $\mathcal{A}_{d}(N_{*})$ defined therein. In the absence of such knowledge, one could empirically estimate the parameters by fitting the decay of $\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}$ on a small test set to the inverse weights $r^{-1}_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})$. A full investigation of this adaptive approach is beyond the scope of this paper, so we reserve it for future work.
A second research direction involves generalizing our method to general Sobolev spaces, thereby removing the periodic boundary conditions inherent to Korobov spaces. A straightforward generalization replaces trigonometric series with finite Walsh series and lattice rules with polynomial lattice rules [10]. While this allows for efficient computation via the Fast Walsh–Hadamard Transform, the resulting approximants are generally discontinuous and lie outside the target Sobolev space. Consequently, they exhibit visual artifacts compared to smoother alternatives like kernel interpolation. A central open question is whether our strategy can be adapted to produce smooth approximants, a challenge we leave for future study.